OpenStudy (shaylasha):

What is the value of x in the equation 4(2x + 10) = 0?

OpenStudy (iluvsoccer):

Our first step is to distribute.

OpenStudy (iluvsoccer):

4 x 2x + 4 x 10 = 0

OpenStudy (iluvsoccer):

Wait... I got confused a little... @brittanyrolfe, may you check this out? :)

OpenStudy (shaylasha):

yea but when i worked it out i got 5 i just need somemone to check and see if its right

OpenStudy (iluvsoccer):

Okay. To check if it's correct, just replace 5 with x. 4 (2(5) + 10) = 0 First we do 2 x 5 = 10 Then we have 10 + 10 = 20 Then we have 4 + 20 and does that equal 0?

OpenStudy (mathsucks321):

Not 5

OpenStudy (iluvsoccer):

^

OpenStudy (shaylasha):

oh

OpenStudy (iluvsoccer):

May someone take over this one? I forgot how to solve this again... @brittanyrolfe

OpenStudy (mathsucks321):

It is -5

OpenStudy (shaylasha):

yea

OpenStudy (shaylasha):

thanks

OpenStudy (mathsucks321):

4(2x+10)=0 2*-5=-10 -10+10=0 0*4=0

OpenStudy (mathsucks321):

Yw ^.^